The average age of three girls is seventeen years four months. The average age of two of them is seventeen years three months. How old is the third girl?
Found 2 solutions by Boreal, Theo:
Answer by Boreal(15235)

total=average*n, by rewriting the definition of an average
total=17 (1/3), for 4 months is 1/3 of a year *3=52
two have a total of 17 (1/4)*2=34 (1/2)
The third girl is 52-34 (1/2)=17 (1/2) or 17 years 6 months

Answer by Theo(13342)

average age of all three girls is 17 years and 4 months.

average age of two of the girls is 17 years and 3 months.

what is the age of the third girl?

if the average age of all three girls is 17 years and 4 months, then the total age of those three girls must be 3 * (17 years and 4 months) = 51 years and 12 months.

if the average age of two of those girls is 17 years and 3 months, then the total age of those two girls must be 2 * (17 years and 3 months) = 34 years and 6 months.

when you subtract the total age of all three of the girls from the total age of two of those girls, you get the age of the third girl.

51 years and 12 months minus 34 years and 6 months results in 17 years and 6 months.

that's the age of the third girl.